OpenStudy (curiouslocal):

last one: secA/cosA - tanA/cotA =1

OpenStudy (zpupster):

\[\frac{ 1/cosA }{ cosA}=\frac{ 1 }{ \cos^2A } = \sec^2A\] and \[\frac{ tanA }{ 1/tanA } = \tan^2A\]

OpenStudy (curiouslocal):

why did it become 1/cos^2A?

OpenStudy (curiouslocal):

how i mean

OpenStudy (zpupster):

sec = 1/cos

OpenStudy (zpupster):

1/cosA*1/cosA = sec^2A ok??